01. Optima Company is a high-technology organization that produces a mass-storage system. The design of Optima's...

01.

Optima Company is a high-technology organization that produces a mass-storage system. The design of Optima's system is unique and represents a breakthrough in the industry. The units Optima produces combine positive features of both compact and hard disks. The company is completing its fifth year of operations and is preparing to build its master budget for the coming year (20X1). The budget will detail each quarter's activity and the activity for the year in total. The master budget will be based on the following information:

  1. Fourth-quarter sales for 20X0 are 55,000 units.
  2. Unit sales by quarter (for 20X1) are projected as follows:
    First quarter 65,000
    Second quarter 70,000
    Third quarter 75,000
    Fourth quarter 90,000
    The selling price is $400 per unit. All sales are credit sales. Optima collects 85% of all sales within the quarter in which they are realized; the other 15% is collected in the following quarter. There are no bad debts.
  3. There is no beginning inventory of finished goods. Optima is planning the following ending finished goods inventories for each quarter:
    First quarter 13,000 units
    Second quarter 15,000 units
    Third quarter 20,000 units
    Fourth quarter 10,000 units
  4. Each mass-storage unit uses 5 hours of direct labor and three units of direct materials. Laborers are paid $10 per hour, and one unit of direct materials costs $80.
  5. There are 65,700 units of direct materials in beginning inventory as of January 1, 20X1. At the end of each quarter, Optima plans to have 30% of the direct materials needed for next quarter's unit sales. Optima will end the year with the same amount of direct materials found in this year's beginning inventory.
  6. Optima buys direct materials on account. Half of the purchases are paid for in the quarter of acquisition, and the remaining half are paid for in the following quarter. Wages and salaries are paid on the 15th and 30th of each month.
  7. Fixed overhead totals $1 million each quarter. Of this total, $350,000 represents depreciation. All other fixed expenses are paid for in cash in the quarter incurred. The fixed overhead rate is computed by dividing the year's total fixed overhead by the year's budgeted production in units.
  8. Variable overhead is budgeted at $6 per direct labor hour. All variable overhead expenses are paid for in the quarter incurred.
  9. Fixed selling and administrative expenses total $250,000 per quarter, including $50,000 depreciation.
  10. Variable selling and administrative expenses are budgeted at $10 per unit sold. All selling and administrative expenses are paid for in the quarter incurred.
  11. The balance sheet as of December 31, 20X0, is as follows:
    Assets
    Cash $ 250,000
    Direct materials inventory 5,256,000
    Accounts receivable 3,300,000
    Plant and equipment, net 33,500,000
         Total assets $42,306,000
    Liabilities and Stockholders’ Equity
    Accounts payable $ 7,248,000*
    Capital stock 27,000,000
    Retained earnings 8,058,000
         Total liabilities and stockholders’ equity $42,306,000
    * For purchase of direct materials only.
  12. Optima will pay quarterly dividends of $300,000. At the end of the fourth quarter, $2 million of equipment will be purchased.

Required:

Prepare a master budget for Optima Company for each quarter of 20X1 and for the year in total. The following component budgets must be included:

1. Sales Budget (units and budgeted sales in thousands)

Optima Company
Sales Budget
For the Year Ending December 31, 20X1
Qtr. 1 Qtr. 2 Qtr. 3 Qtr. 4 Total
Units
Unit price $ $ $ $ $
Total sales $ $ $ $ $
Answer #1
Optima Company
Sales Budget Calculation
Req 1 Qtr. 1 Qtr. 2 Qtr. 3 Qtr. 4 Total Qtr. 1 Qtr. 2 Qtr. 3 Qtr. 4
Units            65,000            70,000            75,000            90,000            300,000
Unit price                  400                  400                  400                  400                    400
Total sales    26,000,000    28,000,000    30,000,000    36,000,000    120,000,000 =65,000*400 =70,000*400 =75,000*400 =90,000*400